USE PYTHON: 4) [30pts] Write program which finds the followings for the given values between 150 to 1000 a) How many odd numbers are between 150 and 1000 (include 150 and 1000) b) Find the sum of odd numbers which are between 150 and 1000 (include 150 and 1000) c) Find the multiplication of odd numbers which are between 150 and 1000 (include 150 and 1000) d) Find the sum of square of odd numbers which are between 150 and 1000 (include 150 and 1000) 5) 1 10pts! Write a program so that for given M. it computes the sum s (Hint: Use the while-loop . As we see in in the class, you can update s as s-stk and you need to increase the k in each step such that k-k+1) 6) [30ptsl The following table contains earthquake magnitude ranges on the Richter scale and their descriptors: Magnitude Less than 2.0 2.0 to less than 3.0 3.0 to less than 4.0 4.0 to less than 5.0 Descriptor Micro Very minor nor Light
Answer #1

In problem 6 please make sure that you write print statements on a single line.

Solution for Problem 4:

count=0
for i in range(150,1001):
    if i%2==1:
        count=count+1
print('Number of odd numbers are',count)
sum=0
for i in range(150,1001):
    if i%2==1:
        sum=sum+i
print('Sum of odd numbers is',sum)
mul=1
for i in range(150,1001):
    if i%2==1:
        mul=mul*i
print('Multiplication of odd numbers is',mul)
square_sum=0
for i in range(150,1001):
    if i%2==1:
        square_sum=square_sum+i*i
print('Sum of squares of odd numbers is',square_sum)

Solution for Problem 5:

M=int(input("Enter value of M:"))
k=1
s=0
while k<M:
   s=s+1/k
   k=k+1
print('Answer is',s)

Solution for 6:

descriptor=''
magnitude=float(input('Enter magnitude of the earthquake:'))
if magnitude<2.0:
    descriptor='Micro'
    print('magnitude',magnitude,'earthquake is considered to be a ',descriptor,' earthquake')
elif magnitude<3.0:
    descriptor='Very minor'
    print('magnitude',magnitude,'earthquake is considered to be a ',descriptor,' earthquake')
elif magnitude<4.0:
    descriptor='Minor'
    print('magnitude',magnitude,'earthquake is considered to be a ',descriptor,' earthquake')
elif magnitude<5.0:
    descriptor='Light'
    print('magnitude',magnitude,'earthquake is considered to be a ',descriptor,' earthquake')
elif magnitude<6.0:
    descriptor='Moderate'
    print('magnitude',magnitude,'earthquake is considered to be a ',descriptor,' earthquake')
elif magnitude<7.0:
    descriptor='Strong'
    print('magnitude',magnitude,'earthquake is considered to be a ',descriptor,' earthquake')
elif magnitude<8.0:
    descriptor='Major'
    print('magnitude',magnitude,'earthquake is considered to be a ',descriptor,' earthquake')
elif magnitude<10.0:
    descriptor='Great'
    print('magnitude',magnitude,'earthquake is considered to be a ',descriptor,' earthquake')
elif magnitude>=10.0:
    descriptor='Meteoric'
    print('magnitude',magnitude,'earthquake is considered to be a ',descriptor,' earthquake')

Output for 4:
E Python |CAUserstclass PycharmProjectsPython]- . HomeworkLib1.py [Python] - PyCharm Eile Edit Yiew Navigate Code Help Chegg HomeworkLib1 Number of odd numbers are 425 Sum of odd nuxbers is 244375 Multiplication of odd numbers 13 1652744965661487554489086430198606359844105675458070498019071302118079254568666866958334967360595724821925942085476342362597586441326794283066887370 Sum of squares of odd numbers is 166104025 Process finished with exit code 0 4 RunPython Console PEP & blank line at end of file 9:16 PM O Type here to searchOutput for 5:

































Output for 6: